# Computation of attractor dimension and maximal sums of Lyapunov exponents using polynomial optimization

### v1

@@ 2.4.2 Shifted spectrum
Next we give the SOS formulations based on the shifted spectrum approach. Proposition 2.17 below is an SOS relaxation of proposition 2.11 for bounding LE sums. Computations giving such bounds also imply bounds on Lyapunov dimension via proposition 2.7, but potentially sharper bounds on dimension can be computed directly using proposition 2.18, which is an SOS relaxation of proposition 2.12. Unlike our other methods, we do not prove that proposition 2.18 gives sharp results in general.
@@
Let $\frac{\mathrm{d}}{\mathrm{d}t}x(t)=f(x(t))$ with $f\in\mathbb{R}^{n}[x]$, for which a semialgebraic set $\mathcal{B}$ is forward invariant. Let $m^{B}_{k}(x,w^{k})$ be defined as in 56. Let $\Sigma^{\Omega}$ be the quadratic module 77 for the set $\Omega=\mathcal{B}\times\mathbb{R}^{n_{k}}$ whose semialgebraic specification includes the same constraints 81 defining $\mathcal{B}$.
@@
    For each $k\in\{1,\ldots,n\}$, and each pair of polynomial degrees $(\nu,\nu^{\prime})$, the maximum leading LE sum $M_{k}$ among trajectories in $\mathcal{B}$ is bounded above by
@@
    Suppose that the invariant set $\mathcal{B}$ is compact, and that its semialgebraic specification is Archimedean. Then, for each $k$,
@@
    where the infimum in 94 is over $V(x,w^{k})=(w^{k})^{\mathsf{T}}U(x)w^{k}$ with $U\in\mathbb{R}^{n_{k}\times n_{k}}[x]$ a symmetric polynomial matrix, and where $\sigma_{i}(x,w^{k})$ and $\rho_{j}(x,w^{k})$ in the SOS representation 77 are likewise quadratic in $w^{k}$.
@@
The first part follows from the upper bound 64 on $M_{k}$ because the SOS constraints in 94 imply the nonnegativity constraints in 64 from proposition 2.11. For the second part, let $B_{1}$ and $B_{2}$ be arbitrary with $B_{1}>B_{2}>\sup_{x_{0}\in\mathcal{B}}M_{k}(x_{0})$. It suffices to show that there exists a polynomial matrix $U(x)$ such that $V(x,w^{k})=(w^{k})^{\mathsf{T}}U(x)w^{k}$ satisfies the SOS constraints of 94 for $B=B_{1}$ with no restrictions on polynomial degree. Since $B_{2}>\sup_{x_{0}\in\mathcal{B}}M_{k}(x_{0})$, the second part of proposition 2.11 guarantees the existence of a symmetric $U\in C^{1}(\mathbb{R}^{n},\mathbb{R}^{n_{k}\times n_{k}})$ for which $V$ satisfies the pointwise inequalities in 64 with $B=B_{2}$. We will show that this implies existence of the desired polynomial $U$ under the present assumptions.
The pointwise inequalities in 64, with $B=B_{2}$ in the $V=(w^{k})^{\mathsf{T}}Uw^{k}$ case, can be written in the form $(w^{k})^{\mathsf{T}}\hat{Z}_{i}(x)w^{k}\geq 0$, where
$\hat{Z}_{1}(x)$ $=U(x)-I_{n_{k}},$ (96)
$\hat{Z}_{2}(x)$ $=-f(x)\cdot D_{x}U(x)+2\,U(x)(B_{2}\,I_{n_{k}}-Df^{[k]}(x)),$ (97)
with $I_{n_{k}}$ denoting the $n_{k}\times n_{k}$ identity matrix. The pointwise inequalities on $\mathcal{B}\times\mathbb{R}^{n_{k}}$ are equivalent to the condition that both $\hat{Z}_{i}(x)\succeq 0$ for all $x\in\mathcal{B}$. In fact, by scaling $U$ we can satisfy both of these equations with the additional restriction that $\hat{Z}_{1}\succ 0$ strictly on $\mathcal{B}$. Let
$Z_{1}(x)$ $=U(x)-I_{n_{k}},$ (98)
$Z_{2}(x)$ $=-f(x)\cdot D_{x}U(x)+2\,U(x)(B_{1}\,I_{n_{k}}-Df^{[k]}(x)),$ (99)
which are the same definitions with $B_{2}$ replaced by $B_{1}$. Since $B_{1}>B_{2}$, we now have both $Z_{i}(x)\succ 0$ strictly for all $x\in\mathcal{B}$.
Because there exists $U\in C^{1}$ such that both $Z_{i}(x)\succ 0$ on $\mathcal{B}$, and $\mathcal{B}$ is compact, there also exists a polynomial matrix $U\in\mathbb{R}^{n_{k}\times n_{k}}[x]$ for which both $Z_{i}(x)\succ 0$ on $\mathcal{B}$. This follows from a strengthened version of the Stone-Weierstrass approximation theorem [25]. This polynomial $U$ satisfies the SOS constraints of 94 due to $\mathcal{B}$ being Archimedean and $Z_{i}$ being strictly positive definite on $\mathcal{B}$. Recalling that $\mathcal{B}$ is the set where all $g_{i}(x)\geq 0$ and $h_{j}(x)=0$, applying a matrix Positivstellensatz [44, theorem 2] guarantees existence of polynomial matrices $S_{i}(x)$ and $T_{i}(x)$, with the $S_{i}(x)$ being squares of other polynomial matrices, such that
@@
and likewise for $Z_{2}$. The existence of such representations implies that both $(w^{k})^{\mathsf{T}}Z_{i}w^{k}$ belong to the quadratic module $\Sigma^{\Omega}$. Therefore $V=(w^{k})^{\mathsf{T}}Uw^{k}$ satisfies the constraints of 94 with $B=B_{1}$, and the second part of the proposition is proven. ∎
@@
Let $\frac{\mathrm{d}}{\mathrm{d}t}x(t)=f(x(t))$ with $f\in\mathbb{R}^{n}[x]$, for which a semialgebraic set $\mathcal{B}$ is forward invariant on $\mathcal{B}\subset\mathbb{R}^{n}$ with each trajectory $x(t)$ bounded for $t\geq 0$. Let the positive integer $k\leq n$ be such that $\sup_{x_{0}\in\mathcal{B}}M_{k+1}(x_{0})<0$. Let $\Sigma^{\Omega}$ be the quadratic module 77 for the set $\Omega=\mathcal{B}\times\mathbb{R}^{n_{k+1}}\times\mathbb{R}^{n_{k}}$ whose semialgebraic specification includes the polynomial constraints on $x$ in 81. Let each trajectory $x(t)$ in $\mathcal{B}$ be bounded for $t\geq 0$. For each $k\in\{1,\ldots,n\}$, and each pair of polynomial degrees $(\nu,\nu^{\prime})$, the the maximum Lyapunov dimension $d_{L}$ among trajectories in $\mathcal{B}$ is bounded above by
@@
$C_{4}(k,\nu,\nu^{\prime})=\inf_{\begin{subarray}{c}B\in[0,1]\\ V\in\mathbb{R}[x,y^{k+1},w]_{\nu}\end{subarray}}B\quad\text{s.t.}\quad\begin{array}[t]{r}V-\left|w\right|^{2}\in\Sigma^{\Omega}_{\nu^{\prime}}\phantom{,}\\ (1-B)\left|y^{k+1}\right|^{2}S_{k}^{B}\in\Sigma^{\Omega}_{\nu^{\prime}}.\end{array}$ (102)
@@ 2.5 Symmetry exploitation
If an ODE being studied using auxiliary function methods has symmetry, then in many cases the same symmetry can be imposed on the auxiliary functions, leading to symmetries in SOS programs that can be exploited computationally by block diagonalizing the corresponding semidefinite programs (SDPs) [11]. More precisely, an ODE system 1 with right-hand side $F(X)$ is invariant under a map $\Lambda:\mathbb{R}^{n}\to\mathbb{R}^{n}$ if and only if $F$ is $\Lambda$-equivariant, meaning that $F(\Lambda X)=\Lambda F(X)$ for all $X$. In such cases, $X(t)$ solves the ODE if and only if $\Lambda X(t)$ does. The trajectory itself may be $\Lambda$-invariant, meaning that $X(t)=\Lambda X(t)$, or it may not be. For any ODE, the set of such transformations forms a group. We focus on the common situation where each $\Lambda$ is an orthogonal linear transformation on $\mathbb{R}^{n}$, meaning the group of such $\Lambda$ is a subgroup of the orthogonal group $O(n)$. (In some contexts the set of $\Lambda$ would be called the linear representation of a group, rather than a group itself, but we do not draw this distinction.) In many formulations of auxiliary function methods for ODEs, if the ODE is invariant under $\Lambda$ then the results are provably unchanged if the optimization is over the restricted set of $V$ which are $\Lambda$-invariant, meaning $V(X)=V(\Lambda X)$ for all $X$. In the context of bounding time averages, see the appendix of [34] for a proof that $V(X)$ can be taken as $\Lambda$-invariant under certain conditions. For our shifted spectrum methods, the Lyapunov function in B.1 can be taken as $\Lambda$-invariant also. We omit the proof, which closely follows the argument in [34].
In order to apply such results in the context of bounding LE sums and Lyapunov dimension, we must determine the symmetries of ODEs on the augmented state spaces for $(x,z^{k})$ and $(x,w^{k})$, governed by 29 and 55. First, regardless of whether the $x$ ODE has any symmetry, the augmented ODEs have a sign symmetry due to linearity of the $y^{k}$ dynamics. These sign symmetries are summarized by proposition 2.19. Second, if the right-hand side $f(x)$ of the $x$ ODE is equivariant under an orthogonal transformation, this induces equivariance on the right-hand sides of the augmented ODEs. This is summarized by proposition 2.20, which invokes the $k^{th}$ multiplicative compound $\Lambda^{(k)}$ of $\Lambda$ (cf. appendix A). Note that $I_{n}$ denotes the $n\times n$ identity matrix, and $\operatorname*{diag}(A,\dots,B)$ is a block diagonal matrix with square blocks $A,\dots,B$.
@@
    The ODE system 29 is invariant under $\operatorname*{diag}(I_{n},-I_{n_{k}})$.
@@
    The ODE system 37 is invariant under $\operatorname*{diag}(I_{n},\pm I_{n_{k}},\mp I_{n_{k+1}})$.
@@
    The ODE system 55 is invariant under $\operatorname*{diag}(I_{n},-I_{n_{k}})$.
@@
    The ODE system 65 is invariant under $\operatorname*{diag}(I_{n},\pm I_{n_{k+1}},\mp I_{n_{k}})$.
@@
All four parts have very similar proofs, so we prove only the last. The claim amounts to the right-hand side of 65 being equivariant under negation of $y^{k+1}$. For $f(x)$ this statement is trivial. The requirement that $Df^{[k+1]}(x)(-y^{k+1})=-Df^{[k+1]}(x)y^{k+1}$ is also immediate. The final requirement that $S_{k}^{B}(x,-y^{k+1},w)=S_{k}^{B}(x,y^{k+1},w)$ holds because $y^{k+1}$ appears quadratically in both the numerator and denominator in 66. ∎
@@
Let $\mathcal{B}\subset\mathbb{R}^{n}$ be invariant under $\Lambda\in O(n)$. If $\frac{\mathrm{d}}{\mathrm{d}t}x(t)=f(x(t))$ with $f:\mathcal{B}\to\mathbb{R}^{n}$ is invariant under $\Lambda$, then the augmented ODEs are invariant under the following orthogonal transformations.
@@
    The ODE system 65 is invariant under $\operatorname*{diag}(\Lambda,\Lambda^{(k+1)},\Lambda^{(k)})$.
@@
By assumption $f(x)$ is equivariant under the orthogonal transformation $\Lambda$, meaning $f(\Lambda x)=\Lambda f(x)$. The multiplicative compound $\Lambda^{(k)}$ is orthogonal by proposition A.2, thus all transformations in the proposition are orthogonal. The claim is proved by showing that the right-hand sides of the augmented ODEs are equivariant under the claimed transformations. All four parts are very similar, so we prove only the first. For this we must show that $\ell_{k}(\Lambda x,\Lambda^{(k)}z^{k})=\Lambda^{(k)}\ell_{k}(x,z^{k})$. It is sufficient to show that
@@
by orthogonality of $\Lambda^{(k)}$. Oeri and Goluskin [34] showed that $Df(\Lambda x)=\Lambda Df(x)\Lambda^{\mathsf{T}}$ . Taking the $k$th additive compound of this equation gives, via proposition A.3,
@@
Orthogonality then gives 103. The proofs of the other parts follow similarly, noting that 103 remains valid with $y^{k}$ or $w^{k}$ in place of $z^{k}$. ∎
Combining both propositions gives any finite symmetry group for the augmented systems. For example, if $\mathcal{B}$ is invariant and $f$ is equivariant under a group generated by $\{\Lambda_{1},\dots,\Lambda_{j}\}$ then the right-hand sides of the augmented systems 29 and 55 are equivariant under the group generated by
$\left\{\operatorname*{diag}(\Lambda_{1},\Lambda_{1}^{(k)}),\dots,\operatorname*{diag}(\Lambda_{j},\Lambda_{j}^{(k)}),\operatorname*{diag}(I_{n},-I_{n_{k}})\right\},$ (106)
and we can restrict to $V$ that are invariant under the same transformations. This can be enforced by representing $V$ in a symmetry-invariant basis. For sign symmetries this can be done with monomial bases, but general rotations require more sophisticated choices of symmetry-invariant polynomial basis. Note that for the shifted spectrum method proposition 2.17 it is sufficient to consider $V$ that are is quadratic in $w^{k}$, which automatically gives invariance of $V$ with respect to the symmetry of proposition 2.19.

### v2

@@ 2.4.2 Shifted spectrum
Now we give the SOS formulations for the shifted spectrum approach. Proposition 2.18 below is an SOS relaxation of proposition 2.11 for bounding LE sums. Computations giving such bounds also imply bounds on Lyapunov dimension via proposition 2.7, but potentially sharper bounds on dimension can be computed directly using proposition 2.19, which is an SOS relaxation of proposition 2.13. We do not prove that proposition 2.19 gives sharp results in general, as we have for our other methods. We also omit its short proof that, as in the first part of proposition 2.18, simply strengthens nonnegativity conditions into SOS conditions.
@@
Let $x(t)$ evolve by the ODE 6, where $f\in\mathbb{R}^{n}[X]$ and $\mathcal{B}\subset\mathbb{R}^{n}$ is a semialgebraic set that is forward invariant. Let $\Sigma^{\Omega}$ be the quadratic module 74 for the set $\Omega=\mathcal{B}\times\mathbb{R}^{C_{n}^{k}}$, using the same constraints 78 that specify $\mathcal{B}$. Let $m^{B}_{k}(x,w^{k})$ be as defined by 56.
@@
    Suppose each trajectory $x(t)$ in $\mathcal{B}$ is bounded for $t\geq 0$. For each $k\in\{1,\ldots,n\}$, and each pair of polynomial degrees $(\nu,\nu^{\prime})$, the maximum leading LE sum $M_{k}$ among trajectories in $\mathcal{B}$ is bounded above by
@@
    Suppose $\mathcal{B}$ is compact and its semialgebraic specification is Archimedean. Then, for each $k$,
@@
    where the infimum in 91 is restricted to $V(x,w^{k})=(w^{k})^{\mathsf{T}}U(x)w^{k}$ with symmetric polynomial matrix $U\in\mathbb{R}^{C_{n}^{k}\times C_{n}^{k}}[x]$, and where $\sigma_{i}(x,w^{k})$ and $\rho_{j}(x,w^{k})$ in the SOS representation 74 are likewise quadratic in $w^{k}$.
@@
The first part follows from the upper bound 65 on $M_{k}$ because the SOS constraints in 91 imply the nonnegativity constraints in 65. For the second part, let $B_{1}$ and $B_{2}$ be arbitrary with $B_{1}>B_{2}>\sup_{x_{0}\in\mathcal{B}}M_{k}(x_{0})$. It suffices to show that there exists a polynomial matrix $U(x)$ such that $V(x,w^{k})=(w^{k})^{\mathsf{T}}U(x)w^{k}$ satisfies the SOS constraints of 91 for $B=B_{1}$ with no restrictions on polynomial degree.
The second part of proposition 2.11 guarantees the existence of a symmetric $U\in C^{1}(\mathbb{R}^{n},\mathbb{R}^{C_{n}^{k}\times C_{n}^{k}})$ for which $V=(w^{k})^{\mathsf{T}}Uw^{k}$ satisfies the pointwise inequalities in 65 with $B=B_{2}$. Equivalently, this $U$ satisfies $(w^{k})^{\mathsf{T}}Z_{i}(x)w^{k}\geq 0$ for all $x\in\mathcal{B}$ and
$Z_{1}(x)$ $=U(x)-I_{C_{n}^{k}},$ (93)
$Z_{2}(x)$ $=-f(x)\cdot D_{x}U(x)+2\,U(x)[B\,I_{C_{n}^{k}}-Df^{[k]}(x)],$ (94)
where $B=B_{2}$ and $I_{C_{n}^{k}}$ denotes the $C_{n}^{k}\times C_{n}^{k}$ identity matrix. This, in turn, is equivalent to both $Z_{i}(x)$ being positive semidefinite for all $x\in\mathcal{B}$. We instead want both $Z_{i}(x)$ to be strictly positive definite on $\mathcal{B}$, which is achieved for $Z_{1}$ by scaling $U$ larger and is then achieved for $Z_{2}$ by taking $B=B_{1}$.
Because there exists $U\in C^{1}$ such that both $Z_{i}(x)$ are positive definite on $\mathcal{B}$, and $\mathcal{B}$ is compact, there also exists a polynomial matrix $U\in\mathbb{R}^{C_{n}^{k}\times C_{n}^{k}}[x]$ for which both $Z_{i}(x)$ are positive definite. This follows from a strengthened version of the Stone-Weierstrass approximation theorem [26]. This polynomial $U$ satisfies the SOS constraints of 91 due to $\mathcal{B}$ being Archimedean and $Z_{i}$ being positive definite on $\mathcal{B}$. Recalling that $\mathcal{B}$ is the set where all $g_{i}(x)\geq 0$ and $h_{j}(x)=0$, applying a matrix Positivstellensatz [47, theorem 2] guarantees existence of polynomial matrices $S_{i}(x)$ and $T_{i}(x)$, with the $S_{i}(x)$ being squares of other polynomial matrices, such that
@@
and likewise for $Z_{2}$. The existence of such representations implies that both $(w^{k})^{\mathsf{T}}Z_{i}w^{k}$ belong to the quadratic module $\Sigma^{\Omega}$. Therefore $V=(w^{k})^{\mathsf{T}}Uw^{k}$ satisfies the constraints of 91 with $B=B_{1}$, and the second part of the proposition is proved. ∎
@@
Let $x(t)$ evolve by the ODE 6, where $f\in\mathbb{R}^{n}[X]$ and $\mathcal{B}\subset\mathbb{R}^{n}$ is a semialgebraic set that is forward invariant. Assume each trajectory $x(t)$ in $\mathcal{B}$ is bounded for $t\geq 0$. Let the positive integer $k<n$ be such that $\sup_{x_{0}\in\mathcal{B}}M_{k+1}(x_{0})<0$. Let $\Sigma^{\Omega}$ be the quadratic module 74 for the set $\Omega=\mathcal{B}\times\mathbb{R}^{C_{n}^{k+1}}\times\mathbb{R}^{C_{n}^{k}}$, using the same constraints 78 that specify $\mathcal{B}$. For each $k\in\{1,\ldots,n\}$, and each pair of polynomial degrees $(\nu,\nu^{\prime})$, the maximum Lyapunov dimension $d_{L}$ among trajectories in $\mathcal{B}$ is bounded above by
@@
$C_{4}(k,\nu,\nu^{\prime})=\inf_{\begin{subarray}{c}B\in[0,1]\\ V\in\mathbb{R}[x,y^{k+1},w]_{\nu}\end{subarray}}B\quad\text{s.t.}\quad\begin{array}[t]{r}V-\left|w\right|^{2}\in\Sigma^{\Omega}_{\nu^{\prime}}\phantom{,}\\ (1-B)\left|y^{k+1}\right|^{2}p_{k}^{B}\in\Sigma^{\Omega}_{\nu^{\prime}},\end{array}$ (97)
with $C_{4}=+\infty$ if the constraint set is empty.
@@ 2.5 Symmetry exploitation
If an ODE being studied using auxiliary function methods has symmetry, then in many cases the same symmetry can be imposed on the auxiliary functions, leading to symmetries in SOS expressions. Symmetries in SOS programs can be exploited computationally by block diagonalizing the corresponding semidefinite programs (SDPs) [12]. An ODE system 1 with right-hand side $F(X)$ is said to be invariant under a map $\Lambda:\mathbb{R}^{n}\to\mathbb{R}^{n}$ if $F$ is $\Lambda$-equivariant, meaning that $F(\Lambda X)=\Lambda F(X)$ for all $X$. In such cases, $X(t)$ solves the ODE if and only if $\Lambda X(t)$ does. For any ODE, the set of such transformations forms a group. We focus on the common situation where each $\Lambda$ is an orthogonal linear transformation on $\mathbb{R}^{n}$, meaning the group of such $\Lambda$ is a subgroup of the orthogonal group $O(n)$. (In some contexts the set of $\Lambda$ would be called the linear representation of a group, rather than a group itself, but we do not draw this distinction.) In many formulations of auxiliary function methods for ODEs, if the ODE is invariant under $\Lambda$ then the results are provably unchanged if the optimization is over the restricted set of $V$ which are $\Lambda$-invariant, meaning $V(X)=V(\Lambda X)$ for all $X$. Such a result in the context of bounding time averages, which is proved in the appendix of [34], justifies imposing symmetries on $V$ in all of our sphere projection methods. An analogous argument holds for the Lyapunov function conditions in B.1, and this justifies imposing symmetries on $V$ in all of our shifted spectrum methods.
The particular symmetries that can be imposed on $V$, and on the polynomials $\sigma_{i}$ and $\rho_{j}$ appearing in Positivstellansätze, follow from the symmetries of ODEs on augmented state space that underlie our various formulations. First, regardless of whether the $x$ ODE has any symmetry, the augmented ODEs have a sign-change symmetry due to linearity of the $y^{k}$ dynamics. These sign-change symmetries are summarized by proposition 2.20. Second, if the right-hand side $f(x)$ of the $x$ ODE is equivariant under an orthogonal transformation $\Lambda$, this induces corresponding equivariance on the right-hand sides of the augmented ODEs, as summarized by proposition 2.21. In the proposition, $\Lambda^{(k)}$ denotes the $k^{\mathrm{th}}$ multiplicative compound of $\Lambda$ (see appendix A), $I_{n}$ denotes the $n\times n$ identity matrix, and $\operatorname*{diag}(A,B,\dots)$ denotes a block diagonal matrix with square blocks $A,B,\ldots$.
@@
    The ODE system 29 is invariant under $\operatorname*{diag}(I_{n},-I_{C_{n}^{k}})$.
@@
    The ODE system 37 is invariant under $\operatorname*{diag}(I_{n},\pm I_{C_{n}^{k}},\mp I_{C_{n}^{k+1}})$.
@@
    The ODE system 55 is invariant under $\operatorname*{diag}(I_{n},-I_{C_{n}^{k}})$.
@@
    The ODE system 70 is invariant under $\operatorname*{diag}(I_{n},\pm I_{C_{n}^{k}},\mp I_{C_{n}^{k+1}})$.
@@
All four parts have very similar proofs, so we prove only the last. The claim amounts to the right-hand side of 70 being equivariant under negation of $y^{k+1}$. For $f(x)$ this statement is trivial. The requirement that $Df^{[k+1]}(x)(-y^{k+1})=-Df^{[k+1]}(x)y^{k+1}$ is also immediate. The final requirement that $p_{k}^{B}(x,w,-y^{k+1})=p_{k}^{B}(x,w,y^{k+1})$ holds because $y^{k+1}$ appears quadratically in both the numerator and denominator in 71. ∎
@@
Let $\mathcal{B}\subset\mathbb{R}^{n}$ be invariant under $\Lambda\in O(n)$. If $\frac{\mathrm{d}}{\mathrm{d}t}x(t)=f(x(t))$ on $\mathcal{B}$ is invariant under $\Lambda$, then the augmented ODEs are invariant under the following orthogonal transformations.
@@
    The ODE system 70 is invariant under $\operatorname*{diag}(\Lambda,\Lambda^{(k)},\Lambda^{(k+1)})$.
@@
By assumption $f(x)$ is equivariant under $\Lambda$, meaning $f(\Lambda x)=\Lambda f(x)$. Orthogonality of $\Lambda$ implies orthogonality of the multiplicative compound $\Lambda^{(k)}$ by proposition A.2, so all transformations in the proposition are orthogonal. The claim is proved by showing that the right-hand sides of the augmented ODEs are equivariant under the claimed transformations. All four parts are very similar, so we prove only the first. For this we must show that $\ell_{k}(\Lambda x,\Lambda^{(k)}z^{k})=\Lambda^{(k)}\ell_{k}(x,z^{k})$. It is sufficient to show that
@@
by orthogonality of $\Lambda^{(k)}$. Since $Df(\Lambda x)=\Lambda Df(x)\Lambda^{\mathsf{T}}$, taking the $k^{\mathrm{th}}$ additive compound of this equation gives, via proposition A.3,
@@
Orthogonality then gives 98. The proofs of the other parts follow similarly, noting that 98 remains valid with $y^{k}$ or $w^{k}$ in place of $z^{k}$. ∎
Combining both propositions gives a finite symmetry group for the augmented systems. For example, if $\mathcal{B}$ is invariant and $f$ is equivariant under a group generated by $\{\Lambda_{1},\dots,\Lambda_{m}\}$ then the right-hand sides of the augmented systems 29 and 55 are equivariant under the group generated by
$\left\{\operatorname*{diag}(\Lambda_{1},\Lambda_{1}^{(k)}),\dots,\operatorname*{diag}(\Lambda_{m},\Lambda_{m}^{(k)}),\operatorname*{diag}(I_{n},-I_{C_{n}^{k}})\right\}.$ (101)
Thus we can restrict to $V$ that are invariant under the same transformations, and likewise for any $\sigma_{i}$ and $\rho_{i}$ appearing in quadratic module representations. The examples of the next section illustrate such symmetry exploitation for particular ODEs.